# Learning Design-Score Manifold to Guide Diffusion Models for Offline Optimization

### v1

@@ Abstract
Offline optimization [6], also named offline model-based optimization, has emerged as a promising solution, enabling design improvement using pre-collected datasets. This approach has proven valuable in molecule generation [7], protein properties [8], and hardware accelerators [9]. Current methods adopt a unidirectional strategy: (i) Training surrogate models to predict scores from designs. These predicted scores are then utilized by various optimizers to identify the optimal designs (forward modeling) [10, 11]; (ii) Generating designs that are conditioned on desired scores through the use of generative models (backward generation). An example of this approach is the employment of diffusion models [12], wherein small amounts of noise are incrementally added to a design sample, and a neural network is trained to reverse this noise-adding procedure. However, forward methods mislead optimizers with overconfident out-of-distribution predictions [6], while backward methods struggle with out-of-distribution generation (OOG) on unseen conditions [13]. These struggles stem from a deeper oversight: existing works operate in isolated design or score spaces, missing the underlying design-score manifold where optimal designs reside.
@@
Figure 1: Overview of the ManGO framework for offline optimization. (a) Illustration of offline optimization on the unknown black-box function (no environment interaction) using an offline dataset. (b) Learning on the design-score manifold with score-augmented training data. (c) Fidelity estimation via unconditional samples generated by ManGO. (d) Bidirectional conditional generation: At the denoising timestep $t$ for sample $i$, it enables both preferred-score and preferred-design guidance, as well as a fidelity-based activation mechanism for inference-time scaling to enhance generation quality (illustrated via the self-supervised importance sampling (self-IS) method). (e) Conceptual illustration of ManGO, which learns on the design-score manifold and enjoys enhanced out-of-distribution generation (OOG) capability. In contrast, learning in the design space faces the out-of-distribution (OOD) issue on unseen conditions. (f) Case study on superconductor’s temperature optimization [14] to demonstrate the superior OOG performance via ManGO versus the design-space approach across varying top data removal.
@@ Results
In this section, we first present the motivation and key idea of ManGO and visualize its manifold and trajectory generation. We then conduct extensive experimental validation on offline SOO and MOO using Design-Bench [6] and Off-MOO-Bench [15]. Finally, we perform systematic ablation studies to analyze the contributions of ManGO’s core components.
Motivation of Learning on Design-Score Manifold
@@
The proposed ManGO framework explicitly learns the design-score manifold and leverages the underlying manifold geometry to co-generate designs and scores. Specifically, ManGO learns a bidirectional mapping between designs and scores: (i) Design-to-Score Prediction: Given any design configuration, ManGO predicts its corresponding score; (ii) Score-to-Design Generation: For any preferred score, ManGO generates its corresponding design. As illustrated in Figure 1.e, this bidirectional mapping provides ManGO with a robust OOG capability to extrapolate beyond training distributions. Let $\hat{\bm{x}}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ denotes a score-augmented design vector at timestep $t$, where $\bm{x}$ and $\bm{y}$ represent the design and its score, respectively. The denoising update can be represented as:
@@
We conduct a controlled experiment to elucidate the advantage of the bidirectional mapping of ManGO on a superconductor [14] task, an 86-D materials design optimization task to maximize a critical temperature, using DDOM [12] as the design-space baseline. In terms of the 128-shot evaluation, Figure 1.f shows that ManGO achieves a consistent score gain of greater than 0.1 over the design-space approach across varying levels of top-data removal (from 70% to 10%). The gain grows to nearly 0.2 when only 10% of the top data are removed. Regarding the 1-shot evaluation, both methods exhibit comparable performance under severe data removal (70%-30%). However, ManGO demonstrates progressively better results as data availability increases. At 10% data removal, ManGO achieves higher scores than the design-space method with 128 shots. These results confirm that: (i) ManGO captures the bidirectional design-score relationships, enabling robustness to OOG challenges; (ii) ManGO exhibits nonlinear scaling of sample efficiency with data quality, achieving superior few-shot performance.
Figure 2: Visualization of manifold and trajectory generation (a-b), unconditional and conditional generation to match preferred scores (c-d). The generative manifold, constructed via ManGO’s design-to-score prediction within the designs’ feasible region, is compared against the original manifold. Generated trajectories visualize ManGO’s score-to-design mapping under specified minimal score and design constraints. Unconditional and conditional samples are generated via ManGO without guidance and with preferred scores as guidance, respectively. (a-b) Manifold and trajectory comparisons for the Branin (SOO) and OmniTest (MOO) tasks. (c) Branin task: Unconditional samples (green) match preferred scores from the training dataset, while conditional samples (blue) surpass the training dataset minimum (grey dashed line). (d) OmniTest task: Conditional samples better approximate preferred scores and Pareto-dominate the training data (grey) compared to unconditional samples.
Figure 3: Pareto front generation under different guidance conditions. Across all subfigures of (a) RE21, (b) ZDT3, (c) DTLZ7, and (d) RE41, columns from left to right respectively show the results of no guidance, standard guidance, and the proposed self-IS-based guidance, demonstrating progressive improvement in generation quality.
Manifold and Trajectory Generation
@@
Table 1: The 100th percentile normalized score ($k=128$) in the Design-Bench benchmark, where the best and runner-up results on each task are bold and underlined numbers. $\mathcal{D}^{\text{(best)}}_{\text{train}}$ denotes the best score in the offline training dataset. Each task’s results are normalized by the best scores in the unobserved dataset.
@@
Evaluation on Single-Objective Optimization
We employ five representative tasks from Design-Bench [6] and sample $10{,}000$ offline design samples per task  [17]: (i) Ant Morphology [18] (60-D parameter optimization for quadruped locomotion speed), (ii) D’Kitty Morphology [19] (56-D parameter optimization for movement efficiency enhancement of a quadruped robot), and (iii) Superconductor [14], (iv) TF-Bind-8 [20] and (v) TF-Bind-10 [20] (discrete DNA sequence optimization for transcription factor binding affinity with sequence lengths 8 and 10). We follow the maximization setting of Design-Bench and normalize scores based on the maximal score in the unobserved dataset [6], where higher scores indicate better performance.
Table 2: Averaged normalized HV and IGD values of synthetic tasks (upper) and RE tasks (lower) in the Off-MOO benchmark, where the best and runner-up results on each task are highlighted by bold and underlined numbers. Higher HV values indicate better performance, while lower IGD values are preferred. $\mathcal{D}^{\text{(best)}}_{\text{train}}$(Preferred) denotes the best (preferred) HV/IGD in the offline training dataset. For compact presentation, reported numbers represent tasks’ performance averaged by the objective number $n$. The sets of ZDT($n$=2), OmniTest($n$=2), and DTLZ($n$=3) consist of 5 ZDT tasks [21], 1 OmniTest task [22], and 2 DTLZ tasks [23], respectively. The sets of RE($n=2$), RE($n=3$), and RE($n=4$) comprise 5, 7, and 2 real-world application tasks [24] with $n=2$, $3$, and $4$, respectively. Note that each task’s results are normalized by the best HV and IGD of its training dataset, and RE($n$=2) presents higher averaged IGD values because the RE22 task has ten times more IGD value than other tasks.
ZDT($n$=2) OmniTest($n$=2) DTLZ($n$=3) Mean Rank
@@
MM-NSGA2($k$=1) 0.888 + 0.013 2.555 + 0.118 1.044 + 0.010 0.454 + 0.129 0.987 + 0.012 0.635 + 0.048 8.3 / 10 6.7 / 10
MO-DDOM ($k$=1) 0.948 + 0.009 1.674 + 0.150 0.983 + 0.002 0.827 + 0.046 0.932 + 0.014 0.795 + 0.063 9.3 / 10 8.0 / 10
ManGO ($k$=1) 1.097 + 0.004 0.729 + 0.064 1.050 + 0.002 0.399 + 0.068 0.971 + 0.016 0.739 + 0.104 6.0 / 10 6.0 / 10
$\text{ManGO}^{\text{+Self-IS}}$ ($k$=1) 1.099 + 0.005 0.702 + 0.065 1.050 + 0.001 0.359 + 0.017 1.053 + 0.015 0.250 + 0.084 4.3 / 10 3.7 / 10
MM-MOBO 0.963 + 0.007 4.723 + 0.164 1.056 + 0.000 0.206 + 0.019 1.075 + 0.000 0.362 + 0.016 4.0 / 10 6.0 / 10
ParetoFlow 1.000 + 0.008 2.867 + 0.405 0.953 + 0.057 1.523 + 0.567 0.998 + 0.009 0.672 + 0.115 7.7 / 10 8.3 / 10
MM-NSGA2 ($k$=256) 1.055 + 0.003 3.592 + 0.044 1.046 + 0.002 1.008 + 0.019 1.086 + 0.000 0.752 + 0.016 4.0 / 10 9.0 / 10
MO-DDOM ($k$=256) 0.981 + 0.006 1.052 + 0.144 1.033 + 0.001 0.270 + 0.010 1.054 + 0.006 0.255 + 0.044 6.7 / 10 4.3 / 10
ManGO ($k$=256) 1.107 + 0.002 0.420 + 0.030 1.051 + 0.002 0.118 + 0.015 1.066 + 0.003 0.172 + 0.013 2.7 / 10 1.7 / 10
$\text{ManGO}^{\text{+Self-IS}}$ ($k$=256) 1.106 + 0.002 0.445 + 0.043 1.052 + 0.000 0.094 + 0.002 1.079 + 0.003 0.123 + 0.033 2.0 / 10 1.3 / 10
RE($n$=2) RE($n$=3) RE($n$=4) Mean Rank
@@
MM-NSGA2($k$=1) 1.016 ± 0.004 56.958 ± 12.159 0.935 ± 0.007 2.979 ± 0.159 0.780 ± 0.009 1.504 ± 0.060 9.3 / 10 9.7 / 10
MO-DDOM ($k$=1) 1.010 ± 0.012 2.261 ± 0.220 1.046 ± 0.002 0.579 ± 0.020 1.058 ± 0.003 0.845 ± 0.008 8.7 / 10 4.7 / 10
ManGO ($k$=1) 1.024 ± 0.004 6.809 ± 0.740 1.051 ± 0.011 0.782 ± 0.083 1.234 ± 0.009 0.421 ± 0.018 5.7 / 10 6.0 / 10
$\text{ManGO}^{\text{+Self-IS}}$ ($k$=1) 1.022 ± 0.004 4.717 ± 2.081 1.066 ± 0.005 0.588 ± 0.027 1.240 ± 0.016 0.304 ± 0.020 4.7 / 10 4.3 / 10
MM-MOBO 1.027 ± 0.002 1.156 ± 0.133 1.071 ± 0.001 0.912 ± 0.093 1.123 ± 0.009 0.816 ± 0.029 3.7 / 10 5.0 / 10
@@
MM-NSGA2 ($k$=256) 1.034 ± 0.000 98.945 ± 0.011 1.069 ± 0.001 2.160 ± 0.029 1.249 ± 0.012 0.320 ± 0.052 2.0 / 10 6.7 / 10
MO-DDOM ($k$=256) 1.020 ± 0.001 1.476 ± 0.031 1.056 ± 0.001 0.552 ± 0.007 1.073 ± 0.004 0.808 ± 0.008 6.7 / 10 3.3 / 10
ManGO ($k$=256) 1.026 ± 0.002 3.373 ± 0.428 1.063 ± 0.005 0.611 ± 0.026 1.248 ± 0.006 0.388 ± 0.014 4.0 / 10 4.7 / 10
$\text{ManGO}^{\text{+Self-IS}}$ ($k$=256) 1.028 ± 0.001 2.264 ± 0.237 1.072 ± 0.002 0.498 ± 0.013 1.264 ± 0.002 0.234 ± 0.020 1.3 / 10 2.0 / 10
Evaluation on Multi-Objective Optimization
We utilize Off-MOO-Bench [15] and sample $60{,}000$ samples per task  [25]: (i) Synthetic Functions (an established collection of MOO evaluation tasks with $2-3$ objectives exhibiting diverse PF characteristics, such as ZDT [21] and DTLZ [23]), and (ii) real-world engineering (RE) applications [24] (a suite of practical design tasks with $2-4$ competing objectives, such as four-bar truss design and rocket injector design). We employ two standard evaluation metrics: (i) Hypervolume (HV) [26], which quantifies the dominated volume between candidate designs and nadir point (each dimension of which corresponds to the worst value of one objective), and (ii) Inverted Generational Distance (IGD) [27], which measures the average minimum distance between candidate designs and the ground-true PF, both metrics applied to non-dominated sorting [28] with $k=256$ candidate designs (100th percentile). While generating high-quality single solutions from purely offline data remains challenging, we also report our method’s performance at $k=1$ to demonstrate competitive performance. Note that while we replace the online query in MOBO/NSGA2 with a surrogate forward model for offline adaptation, performance degradation occurs versus online operation as the surrogate model cannot perfectly emulate environment feedback.
Figure 4: Ablation studies of ManGO’s key components. (a-b) Flexibility to preferred scores on Ant and Superconductor tasks, showing performance versus guided score deviation from true optimum. (c-d) Fidelity threshold analysis: offline SOO (c) and MOO (d) task performance gains versus baseline fidelity (dashed lines mark optimal thresholds). (e-h) Inference-time scaling comparisons on ZDT3 (e-f) and RE21 (g-h): HV (e,g) and IGD (f,h) versus number of function evaluations (NFE) for three approaches, including standard denoising, self-IS scaling, and FKS scaling.
Ablation Study
Robustness to Preferred Scores
@@
Optimal Fidelity Threshold for Inference-Time Scaling
@@
Inference-Time Scaling Beyond More Denoising Steps

### v2

@@ Abstract
Offline optimization [6], also named offline model-based optimization, has emerged as a promising solution, enabling design improvement using pre-collected datasets. This approach has proven valuable in molecule generation [7], protein properties [8], and hardware accelerators [9]. Current methods adopt a unidirectional strategy: (i) Training surrogate models to predict scores from designs. These predicted scores are then utilized by various optimizers to identify the optimal designs (forward modeling) [10, 11]; (ii) Generating designs that are conditioned on desired scores through the use of generative models (backward generation). An example of this approach is the employment of diffusion models [12], wherein small amounts of noise are incrementally added to a design sample, and a neural network is trained to reverse this noise-adding procedure. However, forward methods mislead optimizers with overconfident out-of-distribution predictions [6], while backward methods struggle with out-of-distribution generation (OOG) on unseen conditions [13, 14]. These struggles stem from a deeper oversight: existing works operate in isolated design or score spaces, missing the underlying design-score manifold where optimal designs reside.
@@
Figure 1: Overview of the ManGO framework for offline optimization. (a) Illustration of offline optimization: it identifies optimal designs for an unknown black-box function using an offline dataset (no environment interaction), where designs represent function inputs and scores correspond to outputs. (b) Training a diffusion model on score-augmented data to learn the joint design-score manifold. (c) Fidelity estimation via unconditional samples generated by the trained ManGO model: the fidelity metric determines whether to activate inference-time scaling during conditional generation. (d) Bidirectional conditional generation: it leverages preferred-score or preferred-design conditions to generate corresponding designs or scores, illustrated via the self-supervised importance sampling (self-IS) method at denoising timestep $t$ for sample ii. (e) Conceptual illustration of ManGO: it learns on the design-score manifold to enhance out-of-distribution generation (OOG) capability, contrasted with design-space learning that struggles with OOG issues under unseen conditions [13]. (f) Case study on superconductor’s temperature optimization [15]: it demonstrates the superior OOG performance via ManGO versus the design-space approach (i.e., DDOM) across varying ratios of top data removal.
@@ Results
In this section, we first introduce preliminaries on offline optimization, the basics of diffusion models, baseline methods and performance metrics. Subsequently, to show the motivation and advantages of ManGO, we compare the learned versus the original design-score manifold and visualize the trajectory generation. We then conduct extensive experimental validation on offline SOO and MOO using Design-Bench [6] and Off-MOO-Bench [16]. Finally, we perform systematic ablation studies to analyze the contributions of ManGO’s core components.
Preliminaries
Offline SOO aims to identify the optimal design $\bm{x}^{*}=\arg\min_{\bm{x}\in\mathcal{X}}f(\bm{x})$ using only a pre-collected offline dataset $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$, where $\bm{x}_{i}$ denotes a specific design (also referred to as a solution) and $y_{i}=f(\bm{x}_{i})$ represents its corresponding score (or objective value). Offline MOO aims to identify a set of designs that achieve optimal trade-offs among conflicting objectives using a pre-collected dataset $\mathcal{D}=\left\{(\bm{x}_{i},\bm{y}_{i})\right\}_{i=1}^{N}$, where $\bm{y}_{i}$ denotes the vector of scores corresponding to design $\bm{x}_{i}$. The problem is defined as [17]: $\text{Find }\bm{x}^{*}\in\mathcal{X}\text{ such that }\nexists\bm{x}\in\mathcal{X}\text{ with }\bm{f}(\bm{x})\prec\bm{f}(\bm{x}^{*}),$ where $\prec$ denotes Pareto dominance. A design $\bm{x}^{\prime}$ is said to Pareto dominate another design $\bm{x}$, denoted as $\bm{f}(\bm{x}^{\prime})\prec\bm{f}(\bm{x})$, if $\exists i\in\{1,\ldots,m\},f_{i}(\bm{x}^{\prime})<f_{i}(\bm{x})$ and $\forall j\in\{1,\ldots,m\},f_{j}(\bm{x}^{\prime})\leq f_{j}(\bm{x}).$ Namely, $\bm{x}^{\prime}$ is superior to $\bm{x}$ in at least one objective while being at least as good in all others. A design $\bm{x}^{*}$ is Pareto optimal if no other design $\bm{x}\in\mathcal{X}$ Pareto dominates $\bm{x}^{*}$. The set of all Pareto optimal designs is referred to as the Pareto set (PS), and the set of their scores $\{\bm{f}(\bm{x}^{*})\mid\bm{x}^{*}\in\text{PS}\}$ constitutes the Pareto front. The goal of offline MOO is to identify the PS using a pre-collected dataset, thereby achieving optimal trade-offs among conflicting objectives.
Diffusion models are a type of deep generative models that learn to reverse a gradual noising process, transforming random noise into realistic data through iterative denoising. Let $\bm{x}_{t}$ denote the state of a data sample $\bm{x}_{0}$ at time $t\in[0,T]$, where $\bm{x}_{0}$ is drawn from an unknown data distribution $p_{0}(\bm{x})$. Here, $\bm{x}_{t}$ represents a noisy version of $\bm{x}_{0}$ at time $t$, and $\bm{x}_{T}$ corresponds to a point sampled from a prior noise distribution $p_{T}(\bm{x})$, typically chosen as the standard normal distribution $p_{T}(\bm{x})=\mathcal{N}(\bm{0},\bm{I})$.
The forward diffusion process, also known as the noise-adding process, can be modeled as a stochastic differential equation (SDE) [18]: $d\bm{x}=\mathbf{f}(\bm{x},t)dt+g(t)d\bm{w},$ where $\bm{w}$ denotes the standard Wiener process, $\mathbf{f}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, and $g(t):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient of $\bm{x}_{t}$. The denoising process is defined by the reverse-time SDE: $d\bm{x}=\left[\mathbf{f}(\bm{x},t)-g(t)^{2}\nabla_{\bm{x}}\log p_{t}(\bm{x})\right]dt+g(t)d\tilde{\bm{w}}$, where $dt$ represents an infinitesimal step backward in time, and $d\tilde{\bm{w}}$ is the reverse-time Wiener process.
Baseline Methods and Performance Metrics
We consider existing baseline methods for offline SOO based on three methodological paradigms: (i) Surrogate-based methods: optimizing with surrogate models, including BO-qqEI [19, 20], CMA-ES [10], REINFORCE [21], Gradient Ascent and its variants of mean ensemble and min ensemble. (ii) Forward-modeling methods: employing advanced neural networks like generative models as surrogate models and integrating with surrogate-based methods, including COMs [22], RoMA [23], IOM [24], BDI [11], ICT [25], Tri-Mentoring [26], PGS [27], FGM [28], Match-OPT [29], and RaM [30]. (iii) Inverse-modeling methods: applying score as a condition to reverse design with generative models, including CbAS [31], MINs [32], DDOM [12], BONET [33], and GTG [34].
In offline optimization, where environment interaction is prohibited, it is essential to evaluate multiple candidate solutions; thus, standard benchmarks adopt $k$-shot evaluation with the 100th percentile (best candidate) as the performance metric [6, 16]. All results are normalized using task-specific references for comparison: For Design-Bench with maximization tasks (Table 1), we use standardization normalization for the Superconduct task and min-max normalization for other tasks based on the unobserved dataset’s highest score. For the Off-MOO-Bench with minimization tasks (Tables 2), we use min-max normalization with the best HV and IGD values of training datasets.
Figure 2: Visualization of manifold learning, trajectory generation, and generation capabilities of ManGO. Note that unconditional and conditional samples are generated via ManGO without guidance and with preferred-score guidance, respectively. (a-b) Manifold and trajectory comparisons for the Branin (SOO) and OmniTest (MOO) tasks. The generated manifold is constructed via ManGO’s design-to-score prediction within the feasible region of designs. Close alignment between the ManGO-generated and original manifold, confirming the model’s proficiency in learning complex design-score relationships. Generated trajectories visualize ManGO’s score-to-design mapping under minimal score and design constraints, highlighting its capacity to perform targeted denoising toward desired regions. (c) Branin task: Unconditional samples (green) match preferred scores from the training dataset, while conditional samples (blue) extrapolate beyond the training minimum (grey dashed line). (d) OmniTest task: Conditional samples better approximate preferred scores and Pareto-dominate the training data (grey) compared to unconditional samples. These results indicate that ManGO effectively reconstructs in-distribution samples during unconditional generation—reflecting well-learned manifold structure—while enabling OOG of superior samples through conditional guidance, demonstrating robust extrapolation based on the learned manifold.
Figure 3: Pareto front generation under different guidance conditions. Across all subfigures of (a) RE21, (b) ZDT3, (c) DTLZ7, and (d) RE41, columns from left to right respectively show the results of no guidance, standard guidance, and the proposed self-IS-based guidance. The progressive improvement in generation quality highlights ManGO’s capability in OOG under conditional guidance. Furthermore, the enhanced performance with self-IS-based guidance illustrates the ManGO’s feasibility for more delicate guidance mechanisms.
Motivation and Advantages of ManGO
Compared to conventional manifold learning, our diffusion-based approach provides unique advantages for offline optimization. Unlike conventional methods like kernel-based methods that struggle with complex nonlinear geometries [37], diffusion models excel at capturing intricate manifold structures through stochastic denoising. Crucially, diffusion models inherently support conditional generation, enabling direct generation of high-performing designs conditioned on target scores–a capability absent in standard manifold learning. Furthermore, compared to GANs or VAEs, diffusion models offer superior training stability and generation fidelity [38], which are critical for reliable optimization from offline data. This combination of strong representational capacity and built-in conditional generation makes diffusion models suited for learning the design-score manifold.
@@
Compared to existing methods for offline optimization, our proposed ManGO framework explicitly learns the design-score manifold and leverages the underlying manifold geometry to co-generate designs and scores. Specifically, ManGO learns a bidirectional generation between designs and scores: (i) Design-to-Score Prediction: Given any design configuration, ManGO predicts its corresponding score; (ii) Score-to-Design Generation: For any preferred score, ManGO generates its corresponding design. As illustrated in Figure 1.e, this bidirectional mapping provides ManGO with a robust OOG capability to extrapolate beyond training distributions. Let $\hat{\bm{x}}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ denotes a score-augmented design vector at timestep $t$, where $\bm{x}$ and $\bm{y}$ represent the design and its score, respectively. The denoising update can be represented as:
@@
We conduct a controlled experiment to elucidate the advantage of the bidirectional mapping of ManGO on a superconductor [15] task, an 86-D materials design optimization task to maximize a critical temperature, using DDOM [12] as the design-space baseline. In terms of the 128-shot evaluation, Figure 1.f shows that ManGO achieves a consistent score gain of greater than 0.1 over the design-space approach across varying levels of top-data removal (from 70% to 10%). The gain grows to nearly 0.2 when only 10% of the top data are removed. Regarding the 1-shot evaluation, both methods exhibit comparable performance under severe data removal (from 70% to 30%). However, ManGO demonstrates progressively better results as data availability increases. At 10% data removal, ManGO achieves higher scores than the design-space method with 128 shots. These results confirm that: (i) ManGO captures the bidirectional design-score relationships, enabling robustness to OOG challenges; (ii) ManGO exhibits nonlinear scaling of sample efficiency with data quality, achieving superior few-shot performance.
Manifold and Trajectory Generation Visualization
@@
Evaluation on Single-Objective Optimization
We employ five representative tasks from Design-Bench [6] and sample 10,00010{,}000 offline design samples per task [30]: (i) Ant Morphology [39] (60-D parameter optimization for quadruped locomotion speed), (ii) D’Kitty Morphology [40] (56-D parameter optimization for movement efficiency enhancement of a quadruped robot), and (iii) Superconductor [15], (iv) TF-Bind-8 [41] and (v) TF-Bind-10 [41] (discrete DNA sequence optimization for transcription factor binding affinity with sequence lengths 8 and 10). We follow the maximization setting of Design-Bench and normalize scores based on the maximal score in the unobserved dataset [6], where higher scores indicate better performance.
Evaluation on Multi-Objective Optimization
We utilize Off-MOO-Bench [16] and sample 60,00060{,}000 samples per task [36]: (i) Synthetic Functions (an established collection of MOO evaluation tasks with 232-3 objectives exhibiting diverse PF characteristics, such as ZDT [42] and DTLZ [43]), and (ii) real-world engineering (RE) applications [44] (a suite of practical design tasks with 242-4 competing objectives, such as four-bar truss design and rocket injector design). We employ two standard evaluation metrics: (i) Hypervolume (HV) [45], which quantifies the dominated volume between candidate designs and nadir point (each dimension of which corresponds to the worst value of one objective), and (ii) Inverted Generational Distance (IGD) [46], which measures the average minimum distance between candidate designs and the ground-true PF, both metrics applied to non-dominated sorting [47] with $k=256$ candidate designs (100th percentile). While generating high-quality single solutions from purely offline data remains challenging, we also report our method’s performance at $k=1$ to demonstrate competitive performance. Note that while we replace the online query in MOBO/NSGA2 with a surrogate forward model for offline adaptation, performance degradation occurs versus online operation as the surrogate model cannot perfectly emulate environment feedback.
Table 1: The 100th percentile normalized score ($k=128$) in the Design-Bench benchmark, where the best and runner-up results on each task are bold and underlined numbers. $\mathcal{D}^{\text{(best)}}_{\text{train}}$ denotes the best score in the offline training dataset. Each task’s results are normalized by the best scores in the unobserved dataset. DDOM represents a conditional diffusion-based method.
@@
Table 2: Averaged normalized HV and IGD values of synthetic tasks (upper) and RE tasks (lower) in the Off-MOO benchmark, where the best and runner-up results on each task are highlighted by bold and underlined numbers. Higher HV values indicate better performance, while lower IGD values are preferred. $\mathcal{D}^{\text{(best)}}_{\text{train}}$(Preferred) denotes the best (preferred) HV/IGD in the offline training dataset. For compact presentation, reported numbers represent tasks’ performance averaged by the objective number $n$. The sets of ZDT(nn=2), OmniTest(nn=2), and DTLZ(nn=3) consist of 5 ZDT tasks [42], 1 OmniTest task [48], and 2 DTLZ tasks [43], respectively. The sets of RE($n=2$), RE($n=3$), and RE($n=4$) comprise 5, 7, and 2 real-world application tasks [44] with $n=2$, 33, and $4$, respectively. Note that each task’s results are normalized by the best HV and IGD of its training dataset, and RE(nn=2) presents higher averaged IGD values because the RE22 task has ten times more IGD value than other tasks. Note that MO-DDOM represents a standard conditional diffusion-based baseline method.
ZDT(nn=2) OmniTest(nn=2) DTLZ(nn=3) Mean Rank
@@
$M$-NSGA2(kk=1) 0.888 ± 0.013 2.555 ± 0.118 1.044 ± 0.010 0.454 ± 0.129 0.987 ± 0.012 0.635 ± 0.048 8.3 / 10 6.7 / 10
MO-DDOM (kk=1) 0.948 ± 0.009 1.674 ± 0.150 0.983 ± 0.002 0.827 ± 0.046 0.932 ± 0.014 0.795 ± 0.063 9.3 / 10 8.0 / 10
ManGO (kk=1) 1.097 ± 0.004 0.729 ± 0.064 1.050 ± 0.002 0.399 ± 0.068 0.971 ± 0.016 0.739 ± 0.104 6.0 / 10 6.0 / 10
$\text{ManGO}^{\text{+Self-IS}}$ (kk=1) 1.099 ± 0.005 0.702 ± 0.065 1.050 ± 0.001 0.359 ± 0.017 1.053 ± 0.015 0.250 ± 0.084 4.3 / 10 3.7 / 10
$M$-MOBO 0.963 ± 0.007 4.723 ± 0.164 1.056 ± 0.000 0.206 ± 0.019 1.075 ± 0.000 0.362 ± 0.016 4.0 / 10 6.0 / 10
ParetoFlow 1.000 ± 0.008 2.867 ± 0.405 0.953 ± 0.057 1.523 ± 0.567 0.998 ± 0.009 0.672 ± 0.115 7.7 / 10 8.3 / 10
$M$-NSGA2 (kk=256) 1.055 ± 0.003 3.592 ± 0.044 1.046 ± 0.002 1.008 ± 0.019 1.086 ± 0.000 0.752 ± 0.016 4.0 / 10 9.0 / 10
MO-DDOM (kk=256) 0.981 ± 0.006 1.052 ± 0.144 1.033 ± 0.001 0.270 ± 0.010 1.054 ± 0.006 0.255 ± 0.044 6.7 / 10 4.3 / 10
ManGO (kk=256) 1.107 ± 0.002 0.420 ± 0.030 1.051 ± 0.002 0.118 ± 0.015 1.066 ± 0.003 0.172 ± 0.013 2.7 / 10 1.7 / 10
$\text{ManGO}^{\text{+Self-IS}}$ (kk=256) 1.106 ± 0.002 0.445 ± 0.043 1.052 ± 0.000 0.094 ± 0.002 1.079 ± 0.003 0.123 ± 0.033 2.0 / 10 1.3 / 10
RE(nn=2) RE(nn=3) RE(nn=4) Mean Rank
@@
$M$-NSGA2(kk=1) 1.016 ± 0.004 56.958 ± 12.159 0.935 ± 0.007 2.979 ± 0.159 0.780 ± 0.009 1.504 ± 0.060 9.3 / 10 9.7 / 10
MO-DDOM (kk=1) 1.010 ± 0.012 2.261 ± 0.220 1.046 ± 0.002 0.579 ± 0.020 1.058 ± 0.003 0.845 ± 0.008 8.7 / 10 4.7 / 10
ManGO (kk=1) 1.024 ± 0.004 6.809 ± 0.740 1.051 ± 0.011 0.782 ± 0.083 1.234 ± 0.009 0.421 ± 0.018 5.7 / 10 6.0 / 10
$\text{ManGO}^{\text{+Self-IS}}$ (kk=1) 1.022 ± 0.004 4.717 ± 2.081 1.066 ± 0.005 0.588 ± 0.027 1.240 ± 0.016 0.304 ± 0.020 4.7 / 10 4.3 / 10
$M$-MOBO 1.027 ± 0.002 1.156 ± 0.133 1.071 ± 0.001 0.912 ± 0.093 1.123 ± 0.009 0.816 ± 0.029 3.7 / 10 5.0 / 10
@@
$M$-NSGA2 (kk=256) 1.034 ± 0.000 98.945 ± 0.011 1.069 ± 0.001 2.160 ± 0.029 1.249 ± 0.012 0.320 ± 0.052 2.0 / 10 6.7 / 10
MO-DDOM (kk=256) 1.020 ± 0.001 1.476 ± 0.031 1.056 ± 0.001 0.552 ± 0.007 1.073 ± 0.004 0.808 ± 0.008 6.7 / 10 3.3 / 10
ManGO (kk=256) 1.026 ± 0.002 3.373 ± 0.428 1.063 ± 0.005 0.611 ± 0.026 1.248 ± 0.006 0.388 ± 0.014 4.0 / 10 4.7 / 10
$\text{ManGO}^{\text{+Self-IS}}$ (kk=256) 1.028 ± 0.001 2.264 ± 0.237 1.072 ± 0.002 0.498 ± 0.013 1.264 ± 0.002 0.234 ± 0.020 1.3 / 10 2.0 / 10
Figure 4: Comprehensive ablation studies validating key components of the ManGO framework. (a-b) Flexibility to guidance specification: Performance sensitivity to deviations between guided scores and true optima on Ant (a) and Superconductor (b) tasks, demonstrating ManGO’s stability under suboptimal guidance conditions. (c-d) Fidelity-adaptive scaling: Performance gains relative to baseline fidelity thresholds in offline SOO (c) and MOO (d) tasks, with dashed lines indicating empirically optimal thresholds ($\tau_{\text{opt}}=0.827$ for SOO; $\tau_{\text{opt}}=0.87$ for MOO). (e-h) Inference-time scaling efficiency: HV (e,g) and IGD (f,h) versus number of function evaluations (NFE) for ZDT3 (e-f) and RE21 (g-h) task for three approaches, comparing standard denoising, self-IS scaling, and FKS scaling methods. Consistent performance improvements are achieved through adaptive noise-space exploration.
Ablation Study on Robustness to Preferred Scores
@@
Ablation Study on Optimal Fidelity Threshold
@@
Ablation Study on Inference-Time Scaling